OpenStudy (anonymous):

Simplify square root of negative 121

OpenStudy (anonymous):

11

OpenStudy (anonymous):

\[\sqrt{-121}=\sqrt{121\times -1}=\sqrt{\left( 11 \right)^2\times \iota^2}=11 \iota \]

OpenStudy (redohawk):

-11

OpenStudy (redohawk):

or 11 its all the same!

OpenStudy (redohawk):

if you multiply to negative you will get a positive!

OpenStudy (anonymous):

Wrong @RedoHawk. Try plugging \[\sqrt{-121}\]into a calculator and see what it says.

OpenStudy (marissalovescats):

Welcome to OpenStudy @deleahleath! In this case, technically you cannot take the square root of a negative number, but when we do, we add i because i^2 is equal to -1. So we actually have the square root of 121 *-1 which is the same as square root of 121 * i^2 and 121 squre rooted is 11 and i^2 square rooted is just i. So the square root of -121 is indeed 11i.